Hi, I would like some help with this problem:
In a hospital unit there are eigth nurses and five physicians, seven nurses nd three physicians are females.If a person is selected at random, find the probability that the subject is a nurse or a male.
tried and only get 3/13 and 8/13 but I don't know what to do next.
Thank you

Answer by ewatrrr(24785)

 
Hi,
In a hospital unit there are eigth nurses(7 female) and five physicians(3 female)
One selected at random.
P(nurse 'or' male) = P(nurse) + P(male) - P(male nurse)
= 8/13 + 3/13 - 1/13 = 10/13